Markers Co. has two producing departments: assembly and finishing. The company has been using a plantwide...

  1. Markers Co. has two producing departments: assembly and finishing. The company has been using a plantwide predetermined overhead rate based on direct labor hours. The following estimates were made for the current year:

Assembly

Finishing

Total

Manufacturing Overhead

$200,000

$100,000

$300,000

Direct labor hours

40,000

35,000

75,000

Machine hours

5,000

16,000

21,000

Marker started and completed Job 1512 during the year. The job-order cost sheet indicated the following:

Materials Requisitioned

$18,000

Direct labor cost

16,000

Direct labor hours:

     Assembly

1,700 Hours

     Finishing

1,300 Hours

Machine Hours

     Assembly

1,000 Hours

     Finishing

700 Hours

A total of 2,000 units were produced on Job 1512.

Required:

1.

Assume that Marker uses a plant wide predetermined overhead based on direct labor hours. Calculate the total cost and the unit cost for each of the 2,000 units produced by Job 1512.

2.

Assume that Marker uses separate departmental overhead rates based upon direct labor hours for assembly and machine hours for finishing. Calculate the total cost and the unit cost for each of the 2,000 units produced by Job 1512.
Answer #1

(1) Predetermined OH Rate = Overhead cost/Direct Labor Hours

= $300000/75000 = $4

Total cost & Unit cost :-

Materials $18000
Direct Labor cost $16000
Overhead cost ($4 * 3000) $12000
Total cost $46000
Unit cost ($46000/2000 units) $23

(2) Predetermined OH rate for Assembly = Overhead cost/Direct labor hours

= $200000/40000 = $5

Predetermined OH rate for Finishing = Overhead cost/Machine hours

= $100000/16000 = $6.25

Total cost & Unit cost :-

Materials $18000
Direct Labor cost $16000
Overhead cost :-
Assembly (1700 hours * $5) $8500
Finishing (700 hours * $6.25) $4375
Total cost $46875
Unit cost ($46875/2000 units) $23.44